A body is dropped from a height of h and a second body is thrown vertically up at same instant with a speed of (gh)^1/2 from the ground. They meet at a height.............. from the ground. A. h/ 2 B. h/3 C .2h D. 3h

A  body is dropped from a height of h and a second body is thrown vertically up at same instant with a speed of (gh)^1/2 from the ground. They meet at a height.............. from the ground. 
A. h/ 2
B. h/3
C .2h
D. 3h

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
suppose the balls meet at a distance x from the ground
 
Now,
 
for the ball A
 
h – x = ½ * g* t^2            ..(i)
 
 
for ball B
 
x = \sqrt(gh) t – ½ * g * t^2       …........(ii)
 
Add both the equations
 
h = \sqrt(gh)  t
 
now
 
t = \sqrt(h/g)
 
 
putting this in (i)
 
h – x = h / 2
 
x = h/2
 
hope it helps
 
Regards
Arun (askIITians forum expert)